OpenStudy (anonymous):

HELP PLEASE! METAL. Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -14, and 5 + 8i f(x) = x4 - 362.5x2 + 1450x - 4984 f(x) = x4 - 9x3 + 32x2 - 725x + 4984 f(x) = x4 - 67x2 + 1450x - 4984 f(x) = x4 - 9x3 - 32x2 + 725x - 4984

OpenStudy (anonymous):

\[(x-4)(x+14)\] is part of it

OpenStudy (anonymous):

how do you know johnny? that's my BF.

OpenStudy (anonymous):

then find the quadartic with zeros \(5+8u,5-8i\)

OpenStudy (anonymous):

srry off topic.

OpenStudy (anonymous):

heeeere's johnny!

OpenStudy (anonymous):

Lol.

OpenStudy (anonymous):

the quadratic with zeros \(a+bi\) is \(x^2-2ax+a^2+b^2\)

OpenStudy (anonymous):

so the quadratic with zeros \(5+8i\) is \[x^2-10x+5^2+8^2\]

OpenStudy (anonymous):

your really good at this.

OpenStudy (anonymous):

or \[x^2-10x+89\]

OpenStudy (anonymous):

i can show you why that is true if you like, or you can believe me thank you btw

OpenStudy (anonymous):

that's not an answer choice.

OpenStudy (anonymous):

final job \[(x-4)(x+14)(x^2-10x+89)\] which pretty much is a recipe for failure, use technology

OpenStudy (anonymous):

lol yeah because you were not even close to done!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ugh...

OpenStudy (anonymous):

oh not it is real real easy watch

OpenStudy (anonymous):

tada you can even check that it has the right zeros if you scroll down

OpenStudy (anonymous):

Noo way... your awesome.

OpenStudy (anonymous):

(blush)

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

,,,

OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote of y = 2/9.

OpenStudy (anonymous):

just make the leading coefficient of the numerator 2, the denominator 9 and make sure the degrees are the same you want a simple example you can say \[f(x)=\frac{2x}{9x+1}\]

OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x4 + 16x2

OpenStudy (anonymous):

but you could say \[f(x)=\frac{2x^2-5x+4}{9x^2+3x-2}\] or whatever

OpenStudy (anonymous):

lol what am i the askit casket?

OpenStudy (anonymous):

Your like a genius, I have to ask... :)

OpenStudy (anonymous):

factor out the \(x^2\) and get \[f(x)=x^2(x^2+16)\] then if you want to factor over the complex numbers solve \(x^2+16=0\) get \(x=\pm 4i\) and so \[x\times x\times (x+4i)(x-4i)\] is your linear factorig

OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all potential rational zeros f(x) = x3 - 10x2 + 4x - 24 +/-1, +/-2, +/-3, +/-4, +/-6, +/-8, +/-12, +/-24 +-/-1, +/-One divided by two., +/-2, +/-3, +/-4, +/-6, +/-8, +/-12, +/-24 +/-1, +/-2, +/-3, +/-4, +/-24 +/-1, +/-2, +/-3, +/-4, +/-6, +/-12, +/-24

OpenStudy (anonymous):

this one is for you list ALL factors of 24

OpenStudy (anonymous):

which i think is the first one but you can check

OpenStudy (anonymous):

answer choice 4 then.

OpenStudy (anonymous):

i mean 1

OpenStudy (anonymous):

no answer choice 4 is missing 8

OpenStudy (anonymous):

very tricky

OpenStudy (anonymous):

yea so its 1 right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok, I appreciate you satellite one last question. give me one minute.

OpenStudy (anonymous):

Choice #1: Describe each of the following properties of the graph of the Cosine Function, f(theta) = cos(theta) and relate the property to the unit circle definition of cosine. Amplitude Period Domain Range x-intercepts

OpenStudy (anonymous):

you want the amplitude of cosine?

OpenStudy (anonymous):

that is 1, because cosine is the first coordinate on the unit circle, so goes from -1 to 1

OpenStudy (anonymous):

what's listed... this is like an essay format type thing you know.

OpenStudy (anonymous):

period of cosine is \(2\pi\) as that is the circumference of the unit circle

OpenStudy (anonymous):

domain of cosine is all real numbers, you can take the cosine of anything

OpenStudy (anonymous):

range is from -1 to 1, i.e. \[[-1,1]\]once again because it is a point on the unit circle given by \(a^2+b^2=1\) so the maximum is 1 and the minimum is -1

OpenStudy (anonymous):

it crosses the x axis where the first coordinate on the unit circle is 0, namely at \[\frac{\pi}{2},\frac{3\pi}{2},...\]

OpenStudy (anonymous):

if you want to be fancy you can say \[n\pi-\frac{\pi}{2}\]

OpenStudy (anonymous):

wow, so how are you soo good at this... curious.

OpenStudy (anonymous):

i will let you pad it out with words to look like an essay have fun now i am going to go have a beer with johnny

OpenStudy (anonymous):

if you continue to study math one day you will look at this and it will seem like \(3\times7\)

OpenStudy (anonymous):

I'M A HUGE FAN, BYYE!!

OpenStudy (anonymous):

ttyl

OpenStudy (anonymous):

well I mean i'm not going anywhere, just waiting to see if you are...

OpenStudy (anonymous):

I could do this all day.

OpenStudy (anonymous):

State the horizontal asymptote of the rational function. f(x) = quantity x squared plus nine x minus nine divided by quantity x minus nine. y = 3 None y = 9 y = -9